OpenStudy (anonymous):

Find the roots of the polynomial equation. x^2 + 24 = 11x A. 12, 2 B. 3, 8 C. –8, –3 D. –12, –2

OpenStudy (unklerhaukus):

move all the terms to the left hand side, by talking away 11x form both sides

OpenStudy (unklerhaukus):

\[x^2-11x+24=0\] then you see to factorise into this form \[(x+x_1)(x+x_2)=0\] you need to find \(x_1,x_2\) you need \[x_1+x_2=-11\\x_1x_2=24\]

OpenStudy (anonymous):

to get the roots, this means get the 'answers' if you like or to be more precise get the values that give 0 when you substitute them into an equation. for equations with xsquared as the highest power we get 2 answers, for equations with x-cubed we get 3 answers(roots) and so on. to get the roots of the equation we usually use factors to get them. firstly if the equations has an x-squared or higher then everything moves to one side so you have it all =0. this gives you: \[x^2-11x+24=0\] now you get the factors of this and out them both =0 to get your two x values which are your roots. i'll leave you to get the factors then!

OpenStudy (anonymous):

wich one is it

OpenStudy (anonymous):

hello???

OpenStudy (unklerhaukus):

do you know factors of 24?